# Rotatable Antenna Enabled Wireless Communication and Sensing: Opportunities and Challenges

### v2

@@ IV-B Channel Estimation/Sensing
Beyond this baseline compatibility, RA systems offer new opportunities as well as challenges for coverage/accuracy enhancement in estimation and sensing by actively exploiting the spatial DoFs. Specifically, by sequentially or adaptively scanning the dynamic wireless environment, RAs can achieve multi-view estimation or sensing, thereby enabling more accurate multi-path resolution and/or target separation. However, while these multi-view observations obtained over time can effectively enhance estimation or sensing resolution, they also require more complex space-time signal processing techniques. Furthermore, since RA systems scan across multiple directions to acquire more comprehensive information, the overall sensing and estimation performance critically depends on the rotational scanning time and schedule, necessitating joint coordination and design to achieve the desired estimation or sensing performance. In this context, a viable solution is to employ learning-based frameworks, such as neural networks, to capture correlations across boresight directions, thereby enhancing the estimation and sensing performance for high-dimensional observations or under noisy measurement conditions.
IV-C Boresight Direction Optimization
@@
However, boresight direction optimization also presents several unique challenges. One challenge in RA boresight direction design lies in the practically available discrete boresight direction levels of each RA. Instead of exhaustively searching over all possible discrete levels of all RAs, a practical approach is to first relax discrete constraints and solve the problem for continuous boresight direction values, and then quantize the obtained solutions to their nearest values in the discrete set. Alternatively, heuristic algorithms like greedy search can be employed to efficiently explore the discrete solution space with low complexity. On the other hand, the effectiveness of boresight direction optimization heavily relies on the availability of accurate CSI, which is challenging to acquire in practice. To address this issue, CSI-robust alternatives such as hierarchical search can be adopted, where each antenna first performs coarse scanning to identify promising directional regions and then refines its search within such regions. As for the case without explicit CSI, machine learning-based methods, such as reinforcement learning and neural networks, can be exploited. For example, the RA system can learn a mapping from input features (e.g., location, signal strength, past actions) to boresight direction decisions through offline training or gradually improve its boresight adjustment policy over time via online learning.
IV-D Antenna Configuration
From concept to implementation, the configuration of RA plays a decisive role in shaping system performance and hardware efficiency. In particular, the selection of the boresight control method (electronically or mechanically driven) and rotational scale (array-level or antenna-level rotation) directly affects control accuracy, latency, and hardware cost. Besides control mechanisms, the structure design and deployment strategy of RA arrays further present diverse configuration options. Key factors such as array aperture, inter-antenna spacing, and deployment strategies (centralized versus distributed) significantly impact communication and sensing performance. Therefore, RA configuration warrants careful design in light of varying performance objectives and system constraints.
@@ IV-D Antenna Configuration
On the one hand, the optimal RA configuration largely depends on the system tasks. In wireless sensing systems that prioritize spatial resolution, sparse RA arrays are beneficial, as the large array aperture enables finer sensing resolution. Moreover, the increased inter-antenna spacing helps mitigate coupling effects during rotation. In communication systems, however, a sparse array structure may cause severe inter-user interference when users fall into each other’s beam grating lobes. This highlights the need for context-aware evaluation when applying sparse RA arrays in ISAC systems. Moreover, environment-specific strategies also play a critical role in RA configuration. For instance, in size-constrained applications, compact array designs and control methods, i.e., with closer antenna spacing and electronically driven or MEMS-based control, are preferred to facilitate practical integration. Meanwhile, in rich-scattering or sparsely populated environments, distributed RA deployment is favored to enhance coverage performance and provide macro diversity gains to reduce the detrimental shadowing effects.
@@
Figure 3: The received SNR versus the azimuth angle of the user. (A supplementary video is available at https://youtu.be/r4zzDOmQ2ZI.)
V Case Studies
@@
V-A Experimental Results: Single-User Communication
The experiment is carried out in an indoor environment, where one RA-equipped transmitter operates at a carrier frequency of $5.8$ GHz using 16-quadrature amplitude modulation (QAM). The transmit power is set as $5.8$ dBm. Figure 3 illustrates the received signal-to-noise ratio (SNR) at the user versus its azimuth angle, with the zenith angle fixed at $\theta=0^{\circ}$. As the user’s azimuth angle varies from $-\pi/3$ to $\pi/3$, the RA dynamically adjusts its boresight direction to track and align with the user’s direction, thus maintaining a stable received SNR. In contrast, for the fixed-antenna-based transmitter, user movement results in a growing misalignment between the antenna boresight and the user direction, thus leading to a significant degradation in the received SNR. The above experimental results validate the effectiveness of RA for enhancing communication coverage performance.
V-B Simulation Results: Multi-User Communication and Multi-Target Sensing
@@
Figure 4 shows the max-min signal-to-interference-plus-noise ratio (SINR) versus the maximum zenith angle $\theta_{\mathrm{max}}$ of RA in the multi-user communication system. As $\theta_{\mathrm{max}}$ increases, the RA system gains more spatial flexibility to adapt its radiation pattern to the spatial distribution of users, thus boosting the received SINR more significantly. Moreover, the proposed RA system consistently outperforms both baseline schemes across all values of $\theta_{\mathrm{max}}$, achieving up to $2$ dB and $3.2$ dB improvements compared to random boresight and fixed boresight, respectively. Notably, a significant SINR improvement is observed even when $\theta_{\mathrm{max}}\leq\pi/10$, highlighting that substantial performance gains can be achieved with only a small or moderate range of boresight adjustment when properly optimized.
Figure 5 shows the sensing performance of different schemes for multi-target sensing in terms of spatial power spectrum. It is observed that compared to the random and fixed boresight schemes, the proposed RA-enabled sensing achieves much sharper peaks in the power spectrum aligned with the target directions while maintaining significantly lower sidelobe levels. This demonstrates the enhanced angular resolution and sensing accuracy attributed to the increased directional gain provided by the optimized RA boresight directions.
VI Conclusions
@@
Rui Zhang [F] (rzhang@cuhk.edu.cn) is a Professor with The Chinese University of Hong Kong, Shenzhen, China and the National University of Singapore, Singapore.

### v3

@@ V-B Channel Estimation/Sensing
Beyond this baseline compatibility, RA systems offer new opportunities as well as challenges for coverage/accuracy enhancement in estimation and sensing by actively exploiting the spatial DoFs. By sequentially or adaptively scanning the dynamic wireless environment, RAs can achieve multi-view estimation or sensing, thereby enabling more accurate multi-path resolution and/or target separation. However, while these multi-view observations obtained over time can effectively enhance estimation or sensing resolution, they also require more complex space-time signal processing techniques. To address this, it is essential to develop low-overhead estimation schemes that leverage prior geometric knowledge and spatio-temporal correlations in the environment. A promising approach in this context is to adopt learning-based frameworks, such as neural networks, to capture correlations across boresight directions, thereby enhancing the estimation and sensing performance for high-dimensional observations or under noisy measurement conditions.
V-C Boresight Direction Optimization
@@
However, boresight direction optimization also presents several unique challenges. One challenge in RA boresight direction design lies in the practically available discrete boresight direction levels of each RA. Instead of exhaustively searching over all possible discrete levels of RAs, a practical approach is to first relax discrete constraints and solve the problem for continuous boresight direction values, and then quantize the obtained solutions to their nearest values in the discrete set. Alternatively, heuristic algorithms like greedy search can be employed to efficiently explore the discrete solution space with low complexity. On the other hand, the effectiveness of boresight direction optimization heavily relies on the availability of accurate CSI, which is challenging to acquire in practice. To address this issue, CSI-robust alternatives such as hierarchical search can be adopted, where each antenna first performs coarse scanning to identify promising directional regions and then refines its search within such regions. As for the case without explicit CSI, machine learning-based methods, such as reinforcement learning and neural networks, can be exploited.
V-D Antenna Configuration
From concept to implementation, the configuration of RA plays an essential role in determining both the system performance and hardware efficiency. While RA offers a compact and scalable solution to enhance spatial flexibility, configuring it for practical use remains challenging. In particular, key design aspects, such as boresight control methods, rotational scale, array structure, and deployment strategy, directly affect control accuracy, latency, size, and cost, thereby impacting the overall performance and integration complexity of RA systems. These considerations become even more critical when RAs are deployed on spaceborne or aerial platforms, where additional requirements for miniaturization and ruggedization need to be addressed.
@@ V-D Antenna Configuration
Generally, the optimal RA configuration largely depends on the system tasks. In wireless sensing systems that prioritize spatial resolution, sparse RA arrays are often desirable, as their large array aperture enables finer sensing resolution. Moreover, the increased inter-antenna spacing helps mitigate coupling effects during the boresight rotation. In communication systems, however, a sparse array structure may cause severe inter-user interference when users fall into each other’s beam grating lobes. This underscores the importance of context-aware evaluation when applying sparse RA arrays in ISAC systems. Beyond system functionality, environmental factors further influence RA configuration strategies. For instance, in rich-scattering or sparsely populated environments, distributed RA deployment is preferred to enhance coverage performance and provide macro diversity gains to reduce detrimental shadowing effects. In aerial or spaceborne platforms that are typically size, weight, and power-constrained, compact arrays with small antenna size and electronically driven or MEMS-actuated rotation methods are preferred for practical integration. Furthermore, such platforms are often subject to mechanical vibrations, jitters, and thermal variations, which can induce pointing errors and consequently degrade system performance [15]. To mitigate these effects, ruggedization and robust hardware design strategies should be incorporated.
@@
Figure 3: The received SNR versus the azimuth angle of the user. (Demonstration videos are available at https://youtu.be/r4zzDOmQ2ZI and https://youtu.be/L1aMVaGj5rw.)
VI Case Studies
@@
VI-A Experimental Results: Single-User Communication
The experiment is conducted in an indoor environment, where an RA-equipped transmitter operates at a carrier frequency of $5.8$ GHz using 16-quadrature amplitude modulation (QAM). The transmit power is set to $5.8$ dBm, and the data rate is $0.5$ Mbps. Fig. 3 illustrates the received signal-to-noise ratio (SNR) at the user versus its azimuth angle, with the zenith angle fixed at $\theta=0^{\circ}$. As the user’s azimuth angle varies from $\pi/3$ to $\pi/3$, the RA dynamically adjusts its boresight direction to track and align with the user’s direction, thus maintaining a stable received SNR. In contrast, for the fixed-antenna-based transmitter, user movement results in a growing misalignment between the antenna boresight and the user direction, thus leading to a significant degradation in the received SNR. The above experimental results validate the effectiveness of RA for enhancing communication coverage performance.
VI-B Simulation Results: Multi-User Communication and Multi-Target Sensing
@@
Fig. 4 shows the max-min signal-to-interference-plus-noise ratio (SINR) versus the maximum zenith angle $\theta_{\mathrm{max}}$ of RA in the multi-user communication system. As $\theta_{\mathrm{max}}$ increases, the RA system gains more spatial flexibility to adapt its radiation pattern to the spatial distribution of users, thus boosting the received SINR more significantly. Moreover, the proposed RA system consistently outperforms both baseline schemes across all values of $\theta_{\mathrm{max}}$, achieving up to $2$ dB and $3.2$ dB improvements compared to random boresight and fixed boresight, respectively. Notably, a significant SINR improvement is observed even when $\theta_{\mathrm{max}}\leq\pi/10$, highlighting that substantial performance gains can be achieved with only a small or moderate range of boresight adjustment when properly optimized.
Fig. 5 shows the sensing performance of different schemes for multi-target sensing in terms of spatial power spectrum. It is observed that compared to the random and fixed boresight schemes, the proposed RA-enabled sensing achieves much sharper peaks in the power spectrum aligned with the target directions while maintaining significantly lower sidelobe levels. This demonstrates the enhanced angular resolution and sensing accuracy attributed to the increased directional gain provided by the optimized RA boresight directions.
VII Conclusions
@@
Acknowledgment
The work of Beixiong Zheng was supported in part by the National Natural Science Foundation of China under Grant 62571193, Grant 62201214, and Grant 62331022, the Natural Science Foundation of Guangdong Province under Grant 2023A1515011753, the Guangdong program of under Grant 2023QN10X446 and Grant 2023ZT10X148, the GJYC program of Guangzhou under Grant 2024D01J0079 and Grant 2024D03J0006, and the Fundamental Research Funds for the Central Universities under Grant 2024ZYGXZR087. This work of Changsheng You was supported in part by the National Natural Science Foundation of China under Grant 62571227; in part by the Shenzhen Science and Technology Program under Grant JCYJ20240813094212016; and in part by the Program under Grant 2023QN10X152.
@@ Acknowledgment
Rui Zhang [F] (elezhang@nus.edu.sg) is a Provost’s Chair Professor with the Department of Electrical and Computer Engineering of National University of Singapore, Singapore.